# Dynamics of Spontaneous Topic Changes in Next Token Prediction with Self-Attention

### v3

@@ D.3 Simulation in Section 5
D.4 Additional Experiments for Convergence in Mixed-topics
@@
Building upon the convergence experiments in Li et al., 2024b , our work demonstrates that the correlation coefficients $\langle{\bf W}_{ab}(\tau),{\bf W}^{\text{svm}}_{ab}\rangle/\langle\|{\bf W}_{ab}(\tau)\|_{F},\|{\bf W}^{\text{svm}}_{ab}\|_{F}\rangle$ (green lines) in Figure 8, measured with varying $K=\{6,10,14\}$ and $L=\{8,12,16\}$, approach to 1. These results indicate that Theorem 1 extends beyond individual topics to also capture the convergence dynamics of mixed-topics scenarios, albeit with relatively slower convergence. In these experiments, we fix $T_{train}=4$ and $d=16$. Each point represents the average over 5000 randomly generated instances, trained with 8000 iterations. The shaded area around each line represents the $95\%$ confidence interval, computed over 50 epochs.
D.5 Numerical Analysis for Each Scenario in Figure 3
@@
Let $[\mathbb{S}({\bf X}{\bf W}_{a}(\tau){\bar{\bf x}})]_{i}=a_{i}$ and $[\mathbb{S}({\bf X}{\bf W}_{ab}(\tau){\bar{\bf x}})]_{i}=b_{i}$, for $i\in[T]$. Following Equation 9 and Equation 10, we compute $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{x_{i}}$ and $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{x_{i}}$ to get the highest probability SCC and predict the next token for each input sequence.
@@
Topic continuity. In Fig. 9(a), input sequence $\bf X$ consists of four unique tokens: ${\bf e}_{5}$, ${\bf e}_{1}$, ${\bf e}_{3}$, and ${\bf e}_{4}$. Based on $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left), the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$. Since $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.45$ is the largest, $\widehat{\mathcal{G}}_{a}^{(4)}=\{{\bf e}_{5}\}$ and $\hat{y}_{{\bf W}_{a}}={\bf e}_{5}$. In the mixed-topics scenario, ${\bf W}_{ab}$ preserves the attention priority but ${\bf e}_{4}$ and ${\bf e}_{1}$ have the same priority: ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{4}$, with corresponding $b_{i}$ values: $0.40>0.30>0.15=0.15$. Token ${\bf e}_{5}$ is still with the highest probability to be chosen, as $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.40$. Following the Definition 3, ${\bf W}_{ab}$ keeps topic for the the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$.
Ambiguous sequence. Input sequence $\bf X$ in Fig. 9(b) has two unique tokens: ${\bf e}_{1}$ and ${\bf e}_{4}$. The priority order is ${\bf e}_{1}>{\bf e}_{4}$, following $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left). The corresponding values are $a_{1}=a_{3}=0.3$ and $a_{2}=a_{4}=0.2$. Then $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{1}}=2\times 0.30$ and $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.20$. Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ is $\{{\bf e}_{5}\}$ with the highest probability. ${\bf W}_{ab}$ makes ${\bf e}_{4}$ and ${\bf e}_{1}$ with the same priority, as indicated by $\mathcal{G}^{(4)}_{ab}$ in Figure 3 (left). Both ${\bf e}_{1}$ and ${\bf e}_{4}$ are within the highest probability SCC, $\widehat{\mathcal{G}}_{ab}^{(4)}$, due to $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{1}}=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.25$. Although $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$, $\widehat{\mathcal{G}}_{a}^{(4)}\in\widehat{\mathcal{G}}_{ab}^{(4)}$. Therefore, the sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ is ambiguous, based on the Definition 4.
@@
Appendix E Experimental Details in Section 6
In this section, we provide the experimental details in four LLMs: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3. Here, we outline the general procedure used in each model, under identical parameter settings, to generate continuations for each segment of the abstract as follows:
@@ Appendix E Experimental Details in Section 6
    Randomly select 5 different papers from paper $A$, as papers in $\{B_{i}\}_{i=1}^{5}$.
@@
    For the input segment $\bf{X}$, apply RAG to extract top 3 relevant excerpts (chunks) from paper $A$ as the the knowledge $A$, denoted as $\text{Ref}_{A}$. Each chunk is with 800 tokens.
@@
    Similarly, retrieve top 3 relevant excerpts from paper $B_{i}$ as the knowledge $B_{i}$, denoted as $\text{Ref}_{B_{i}}$.
@@
    Combine the knowledge from Topic $A$ and from Topic $B_{i}$ as the knowledge $AB_{i}$ for mixed-Topic, denoted as $\text{Ref}_{AB_{i}}$.
@@
    For the input segment $\bf{X}$, each LLM with the follow prompts, $\text{Prompt}_{A}$ and $\text{Prompt}_{AB_{i}}$, to generate the continuations as $\hat{y}_{\bf{W}_{a}}$ and $\hat{y}_{\bf{W}_{ab}}$, respectively. Notably, the only difference between $\text{Prompt}_{A}$ and $\text{Prompt}_{AB_{i}}$ is the reference excerpts provided $\text{Ref}_{A}$ or $\text{Ref}_{AB_{i}}$. All LLMs were queried with a temperature of 0 to match the greedy decoding in our theoretical framework. The maximum completion length was set to 1000 tokens to ensure that the generated continuations could complete the abstract.
@@
      $\text{Prompt}_{A}$:
      Here are some relevant excerpts from research paper(s) as reference:$\text{Ref}_{A}$. Below is the 1st fragment of an abstract from arXiv paper $A$: $\bf{X}$. Please continue the 2nd fragment of the abstract based on the relevant excerpts without including the given content in the output.
@@
      $\text{Prompt}_{AB_{i}}$:
      Here are some relevant excerpts from research paper(s) as reference:$\text{Ref}_{AB_{i}}$. Below is the 1st fragment of an abstract from arXiv paper $A$: $\bf{X}$. Please continue the 2nd fragment of the abstract based on the relevant excerpts without including the given content in the output.
@@
    Calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across five pairs of paper $A$ and paper $B_{i}$.
E.1 Impact of the Input Length
To investigate the impact of the input length, we vary $T=\{10,30,50,70,90,110,130,150\}$ for every paper as Topic $A$, increasing the length of the input segment $\bf X$, as shown on the x-axis from Figure 5.
E.2 Impact of the Topic Ambiguity
We quantify topic (paper) ambiguity by the average similarity among each paper’s keywords. As arXiv papers do not provide keywords, we use Llama-3.3 to generate four keywords for each paper prior to generating continuations with the LLMs. To investigate the topic ambiguity, we fix the input length with $T=80$ for every paper as Topic $A$ and order papers by the average keywords similarity, as shown on the x-axis of Figure 5. Higher keywords similarity corresponds to lower topic ambiguity.
Appendix F Computational Resources for Experiments
In our simulations based on the single-layer self-attention model, each group of parameter setting requires 7 hours to train two models separately, one for single input topic and one for mixed-topic, followed by 2 additional hours for next-token prediction.
In our experiments on LLMs, we query GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3 through API calls. All experiments were conducted on a standard laptop without specialized hardware. For each LLM, the full process, including selecting relevant excerpts using RAG and generating continuations, requires approximately 20 hours of runtime, with a total of 13 million input tokens and 2.5 million output tokens. The total API usage cost for the experiments is approximately 200 USD.
Appendix G Impact Statement

### v4

@@ C.3 Simulation in Section 5
C.4 Additional experiments for convergence in mixed topics
@@
Building upon the convergence experiments in li2024mechanics, our work demonstrates that the correlation coefficients $\langle{\bf W}_{ab}(\tau),{\bf W}^{\text{svm}}_{ab}\rangle/\langle\|{\bf W}_{ab}(\tau)\|_{F},\|{\bf W}^{\text{svm}}_{ab}\|_{F}\rangle$ (green lines) in Figure 8, measured with varying $K=\{6,10,14\}$ and $L=\{8,12,16\}$, approach to 1. These results indicate that Theorem 1 extends beyond individual topics to also capture the convergence in mixed-topic scenarios, albeit with relatively slower convergence. In these experiments, we fix $T_{train}=4$ and $d=16$. Each point represents the average over 5000 randomly generated instances, trained with 8000 iterations. The shaded area around each line represents the $95\%$ confidence interval, computed over 50 epochs.
C.5 Numerical analysis for each scenario in Figure 3
@@
Let $\mathbb{S}({\bf X}{\bf W}_{a}(\tau){\bar{\bf x}})]_{i}=a_{i}$ and $\mathbb{S}({\bf X}{\bf W}_{ab}(\tau){\bar{\bf x}})]_{i}=b_{i}$, for $i\in[T]$. Following Equation 9 and Equation 10, we compute $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{x_{i}}$ and $Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{x_{i}}$ to get the highest probability SCC and predict the next token for each input sequence.
@@
Topic continuity. In Fig. 9(a), input sequence $\bf X$ consists of four unique tokens: ${\bf e}_{5}$, ${\bf e}_{1}$, ${\bf e}_{3}$, and ${\bf e}_{4}$. Based on $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left), the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$. Since $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.45$ is the largest, $\widehat{\mathcal{G}}_{a}^{(4)}=\{{\bf e}_{5}\}$ and $\hat{y}_{{\bf W}_{a}}={\bf e}_{5}$. In the mixed-topic scenario, ${\bf W}_{ab}$ preserves the attention priority but ${\bf e}_{4}$ and ${\bf e}_{1}$ have the same priority: ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{4}$, with corresponding $b_{i}$ values: $0.40>0.30>0.15=0.15$. Token ${\bf e}_{5}$ is still with the highest probability to be chosen, as $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.40$. Following the Definition 3, ${\bf W}_{ab}$ keeps topic for the the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$.
Ambiguous sequence. Input sequence $\bf X$ in Fig. 9(b) has two unique tokens: ${\bf e}_{1}$ and ${\bf e}_{4}$. The priority order is ${\bf e}_{1}>{\bf e}_{4}$, following $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left). The corresponding values are $a_{1}=a_{3}=0.3$ and $a_{2}=a_{4}=0.2$. Then $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{1}}=2\times 0.30$ and $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.20$. Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ is $\{{\bf e}_{5}\}$ with the highest probability. ${\bf W}_{ab}$ makes ${\bf e}_{4}$ and ${\bf e}_{1}$ with the same priority, as indicated by $\mathcal{G}^{(4)}_{ab}$ in Figure 3 (left). Both ${\bf e}_{1}$ and ${\bf e}_{4}$ are within the highest probability SCC, $\widehat{\mathcal{G}}_{ab}^{(4)}$, due to $Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{1}}=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.25$. Although $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$, $\widehat{\mathcal{G}}_{a}^{(4)}\in\widehat{\mathcal{G}}_{ab}^{(4)}$. Therefore, the sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ is ambiguous, based on the Definition 4.
@@
Appendix D Experimental details in Section 6
In this section, we provide the experimental details in four LLMs: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3. Here, we outline the general procedure used in each model, under identical parameter settings, to generate continuations for each segment of the abstract.
@@ Appendix D Experimental details in Section 6
    Randomly select 5 papers different from paper A, as papers B in $\{\text{B}_{i}\}_{i=1}^{5}$.
@@
    For the input segment $\bf{X}$, apply RAG to extract top 3 relevant excerpts (chunks) from paper A as the the knowledge A, denoted as $\text{Ref}_{\text{A}}$. Each chunk has 800 tokens length.
@@
    Similarly, retrieve top 3 relevant excerpts from paper $\text{B}_{i}$ as the knowledge $\text{B}_{i}$, denoted as $\text{Ref}_{\text{B}_{i}}$.
@@
    Combine the knowledge from Topic A and from Topic $\text{B}_{i}$ as the knowledge $\text{AB}_{i}$ for mixed Topics, denoted as $\text{Ref}_{\text{AB}_{i}}$.
@@
    For the input segment $\bf{X}$, promot each LLM with $\text{Prompt}_{\text{A}}$ and $\text{Prompt}_{\text{AB}_{i}}$, to generate the continuations as $\hat{y}_{\bf{W}_{a}}$ and $\hat{y}_{\bf{W}_{ab}}$, respectively. Notably, the only difference between $\text{Prompt}_{\text{A}}$ and $\text{Prompt}_{\text{AB}_{i}}$ is the reference excerpts provided $\text{Ref}_{\text{A}}$ or $\text{Ref}_{\text{AB}_{i}}$. All LLMs are set with a temperature of 0 to match the greedy decoding in our theoretical framework. The maximum completion length was set to 1000 tokens to ensure that the generated continuations could complete the abstract.
@@
      $\text{Prompt}_{\text{A}}$:
[…]
@@
      $\text{Prompt}_{\text{AB}_{i}}$:
[…]
@@
    Calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across five pairs of paper A and paper $\text{B}_{i}$.
D.1 Impact of input length
To investigate the impact of the input length, we vary $T=\{10,30,50,70,90,110,130,150\}$ for every paper as Topic $A$, increasing the length of the input segment $\bf X$ extracted from the abstract of paper A, as shown on the x-axis from Figure 5.
D.2 Impact of topic ambiguity
We quantify topic (paper) ambiguity by computing the average similarity among each paper’s keywords. Since arXiv papers do not provide keywords, we use Llama-3.3 to generate four keywords for each paper prior to generating continuations with the LLMs. To investigate the topic ambiguity, we fix the input length with $T=80$ for every paper as Topic A and order papers by the average keywords similarity, as shown on the x-axis of Figure 5. A higher keywords similarity corresponds to lower topic ambiguity.
D.3 Additional experiments
D.3.1 Extended input length
To further examine the impact of the input length, we randomly select 50 out of 100 arXiv papers introduced in Sec. 6 and extend the input length with the first $310,460,\ldots,1210$ words from the introduction of each paper A as the input prompt. We start from $T=310$ because shorter introductions provide insufficient contextual information to reliably extract relevant excerpts from the full paper when using RAG. As shown in Figure 10, the average cosine similarity generally increases with the more introduction content from paper A, with the exception of DeepSeek, which exhibits only a marginal improvement.
Figure 10: Extended length of input sequences.
D.3.2 Alternative measure of ambiguity
As an alternative approach to validate our results, we also measure the ambiguity based on the similarity between the keywords of paper A and those of paper B. Following the same setup, we rank the cosine similarity for each paper A by the increasing similarity between its keywords and those of five papers B, where a higher level corresponds to a greater ambiguity. As shown in Figure 11, our results still hold since the cosine similarity remains relatively stable as the ambiguity level increases.
Figure 11: Topic ambiguity measured by the similarity between the keywords of paper A and those of paper B. Higher keywords similarity indicate greater topic ambiguity between the two papers.
Appendix E Computational resources for experiments
In our simulations based on the single-layer self-attention model, each group of parameter setting requires 7 hours to train two models separately, one for single input topic and one for mixed topics, followed by 2 additional hours for next-token prediction.
In our experiments on LLMs, we query GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3 through API calls. All experiments were conducted on a standard laptop without specialized hardware. For each LLM, the full process, including selecting relevant excerpts using RAG and generating continuations, requires approximately 80 hours of runtime, with a total of 30 million input tokens and 5 million output tokens. The total API usage cost for the experiments is approximately 350 USD.
Appendix F Impact statement